OpenStudy (anonymous):

Please Help Will Give Medal and Fan Simplify The Following Expression: 5^-1 ÷ 5^4 A. 125 B. 1/125 C. 3,125 B. 1/3,125

OpenStudy (anonymous):

b

OpenStudy (anonymous):

1/125

OpenStudy (anonymous):

thx

OpenStudy (skullpatrol):

$$\Huge \frac{1}{5} \cdot \frac{1}{5^4}$$

OpenStudy (skullpatrol):

You have TWO choices marked "B"?

OpenStudy (anonymous):

She knew what I meant. No kind of answer choice has A B C B. She used her common sense.

OpenStudy (anonymous):

A. 125 B. 1/125 C. 3,125 D. 1/3,125

OpenStudy (skullpatrol):

Now you use YOUR common sense and calculate the above expression :)

OpenStudy (skullpatrol):

$$\Huge\frac{1}{5 \cdot 5\cdot 5\cdot5\cdot5}$$